# RealCamo: Boosting Real Camouflage Synthesis with Layout Controls and Textual-Visual Guidance

### v1

@@ A.1 Related Work
As the extension of general object detection (GOD) (CVPR2015FCN-GOD) and salient object detection (SOD) (TIP2015SOD), camouflaged object detection (COD) (CVPR2020SINet-COD10K; ICML2025RUN) is proposed to accurately localize and precisely segment camouflaged targets hidden within complex scenes, which is applicable to numerous real-world scenarios, such as concealed defect detection in industry, pest monitoring in agriculture, and medical lesions diagnosis (AIR2024CODsurvey). SINet (CVPR2020SINet-COD10K) pioneered COD using a biologically inspired two-stage search-and-identify framework, followed by a wide range of bio-inspired (CVPR2021PFNet; TPAMI2022SINetV2; CVPR2022ZoomNet; CVPR2022SegMaR; TIP2023FSNet; TIP2024IdeNet) and feature-enhanced (CVPR2021MGL; CVPR2021LSR-NC4K; TIP2022FAPNet; CVPR2023FSPNet; ICLR2024ICEG; TPAMI2024CamoFormer) methods. Subsequent studies incorporated auxiliary cues to complement RGB images, including edges and textures (IJCAI2022BgNet; MIR2023DGNet; CVPR2023FEDER), depth information (ICCV2023PopNet; ACMMM2023DaCOD; ACMMM2024DSAM), and frequency-domain representations (ACMMM2023FRINet; ACMMM2023FPNet; ECCV2024FSEL). Representative works exploit frequency decomposition with boundary refinement (CVPR2023FEDER), depth-induced saliency enhancement (ICCV2023PopNet), and joint frequency–spatial learning (ECCV2024FSEL). Despite steady progress, the development of COD remains fundamentally limited by the small scale of available training datasets, which drives the development of camouflaged image generation tasks.
@@ A.2 Datasets and Evaluation Metrics
Datasets. We evaluate the effectiveness of our synthetic dataset, SynCOD12K, on the COD task following the standard practice of SINet (CVPR2020SINet-COD10K). Experiments are conducted on four widely used benchmark datasets: CHAMELEON (CHAMELEON), CAMO (CVIU2019CAMO), COD10K (CVPR2020SINet-COD10K), and NC4K (CVPR2021LSR-NC4K). The CHAMELEON dataset contains 76 images, while CAMO consists of 1,250 images across 8 classes. COD10K includes 5,066 images spanning ten super-classes, and NC4K serves as the largest testing benchmark with 4,121 well-annotated images. For standard COD training, we follow the common practice by using 1,000 images from CAMO and 3,040 images from COD10K as the training set, which also aligns with the training protocol for camouflaged image generation in (CVPR2024LAKERED). The remaining images from these two datasets, together with all images from CHAMELEON and NC4K, are used for evaluation.
Evaluation Metrics. To thoroughly evaluate the performance, we adopt five standard metrics widely used in COD and SOD tasks: mean absolute error ($M$) (MAE), mean E-measure ($E_{\phi}$) (E-measure), mean F-measure ($F_{\beta}$) (F-measure), weighted F-measure ($F_{\beta}^{w}$) (wF-measure), and structure measure ($S_{\alpha}$) (S-measure). Better performance is indicated by lower values of $M$ and higher values of $E_{\phi}$, $F_{\beta}$, $F_{\beta}^{w}$, and $S_{\alpha}$.
@@ B.1 Diffusion Models
Tex-to-image (T2I) generation methods, represented by Stable Diffusion (CVPR2022LDM), synthesize images conditioned on textual descriptions ($c^{t}$) via a denoising diffusion paradigm (NIPS2020DDPM). These models can operate either in pixel space or latent space, with latent diffusion models (CVPR2022LDM) offering substantially improved computational efficiency.
@@
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (8)
@@
$\mathcal{L}=\mathbb{E}_{x_{0},t,\epsilon\sim\mathcal{N}(0,I)}\left[\parallel\epsilon-\epsilon_{\theta}(x_{t},t,c^{t})\parallel_{2}^{2}\right].$ (9)
@@
$x_{t-1}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t,c^{t})\right)+\sigma_{t}\epsilon^{\prime},$ (10)
@@
$\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t,c^{t})}{\sqrt{\bar{\alpha}_{t}}}.$ (11)
In latent diffusion models (LDMs) (CVPR2022LDM), the above forward and reverse process is conducted on latent embeddings $z=\mathcal{E}_{\mathrm{VAE}}(x)$, as a latent representation of raw RGB pixels $x$ encoded by a VAE (ICLR2014VAE1; ICML2014VAE2). And the denoised latent representation $z_{0}$ is further decoded into a normal RGB image using the VAE decoder, as $x_{0}=\mathcal{D}_{\mathrm{VAE}}(z_{0})$.
@@ B.2 ControlNet
ControlNet (ICCV2023ControlNet) remarkably extends the controllability of large pretrained T2I LDMs by incorporating addition conditions ($c^{f}$) like Canny edges, human pose, or depth map, which provide structural priors for the target generated image.
@@
$y_{\mathrm{c}}=\mathcal{F}(x;\Theta)+\mathcal{Z}(\mathcal{F}(x+\mathcal{Z}(c^{f};\Theta_{\mathrm{Z}1});\Theta_{\mathrm{c}});\Theta_{\mathrm{Z}2})$ (12)
Training Objective. With the introduction of additional conditioning vector $c^{f}$, the overall training objective (Equation 9) is reformulated as:
$\mathcal{L}=\mathbb{E}_{x_{0},t,c^{t},c^{f},\epsilon\sim\mathcal{N}(0.I)}\left[\parallel\epsilon-\epsilon_{\theta}(x_{t},t,c^{t},c^{f})\parallel_{2}^{2}\right].$ (13)
@@
Appendix C User Study
The results are currently being collected and pending further analysis..

### v2

@@ A.1 Related Work
As the extension of general object detection (GOD) (Long et al., 2015) and salient object detection (SOD) (Borji et al., 2015), camouflaged object detection (COD) (Fan et al., 2020; He et al., 2025) is proposed to accurately localize and precisely segment camouflaged targets hidden within complex scenes, which is applicable to numerous real-world scenarios, such as concealed defect detection in industry, pest monitoring in agriculture, and medical lesions diagnosis (Xiao et al., 2024). SINet (Fan et al., 2020) pioneered COD using a biologically inspired two-stage search-and-identify framework, followed by a wide range of bio-inspired (Mei et al., 2021; Fan et al., 2022; Pang et al., 2022; Jia et al., 2022; Song et al., 2023; Guan et al., 2024) and feature-enhanced (Zhai et al., 2021; Lv et al., 2021; Zhou et al., 2022; Huang et al., 2023; He et al., 2024; Yin et al., 2024) methods. Subsequent studies incorporated auxiliary cues to complement RGB images, including edges and textures (Sun et al., 2022; Ji et al., 2023; He et al., 2023a), depth information (Wu et al., 2023b; Wang et al., 2023; Yu et al., 2024b), and frequency-domain representations (Xie et al., 2023; Cong et al., 2023; Sun et al., 2025). Representative works exploit frequency decomposition with boundary refinement (He et al., 2023a), depth-induced saliency enhancement (Wu et al., 2023b), and joint frequency–spatial learning (Sun et al., 2025). Weakly-supervised COD (He et al., 2023b; Chen et al., 2024; Gao et al., 2025), semi-supervised COD (Lai et al., 2024; Fu et al., 2024; Yue et al., 2025), and unsupervised COD (Du et al., 2025a; Yan et al., 2025; Du et al., 2025b) methods are also explored. Despite steady progress, the development of COD remains fundamentally limited by the small scale of available training datasets, which drives the development of camouflaged image generation tasks.
@@ A.2 Datasets and Evaluation Metrics
Datasets. Downstream application on the COD task following the standard practice of SINet (Fan et al., 2020). Experiments are conducted on four widely used benchmark datasets: CHAMELEON (Skurowski et al., 2018), CAMO (Le et al., 2019), COD10K (Fan et al., 2020), and NC4K (Lv et al., 2021). The CHAMELEON dataset contains 76 images, while CAMO consists of 1,250 images across eight classes. COD10K includes 5,066 images spanning ten super-classes, and NC4K serves as the largest testing benchmark with 4,121 well-annotated images. For standard COD training, we follow the common practice by using 1,000 images from CAMO and 3,040 images from COD10K as the training set, which also aligns with the training protocol for camouflaged image generation in (Zhao et al., 2024). The remaining images from these two datasets, together with all images from CHAMELEON and NC4K, are used for evaluation.
Evaluation Metrics. To thoroughly evaluate the performance, we adopt five standard metrics widely used in COD and SOD tasks: mean absolute error ($M$) (Perazzi et al., 2012), mean E-measure ($E_{\phi}$) (Fan et al., 2018), mean F-measure ($F_{\beta}$) (Achanta et al., 2009), weighted F-measure ($F_{\beta}^{w}$) (Margolin et al., 2014), and structure measure ($S_{\alpha}$) (Fan et al., 2017). Better performance is indicated by lower values of $M$ and higher values of $E_{\phi}$, $F_{\beta}$, $F_{\beta}^{w}$, and $S_{\alpha}$.
@@ B.1 Diffusion Models
Tex-to-image (T2I) generation methods, represented by Stable Diffusion (Rombach et al., 2022), synthesize images conditioned on textual descriptions ($c^{t}$) via a denoising diffusion paradigm (Ho et al., 2020). These models can operate either in pixel space or latent space, with latent diffusion models (Rombach et al., 2022) offering substantially improved computational efficiency.
@@
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (S1)
@@
$\mathcal{L}=\mathbb{E}_{x_{0},t,\epsilon\sim\mathcal{N}(0,I)}\left[\parallel\epsilon-\epsilon_{\theta}(x_{t},t,c^{t})\parallel_{2}^{2}\right].$ (S2)
@@
$x_{t-1}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t,c^{t})\right)+\sigma_{t}\epsilon^{\prime},$ (S3)
@@
$\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t,c^{t})}{\sqrt{\bar{\alpha}_{t}}}.$ (S4)
In latent diffusion models (LDMs) (Rombach et al., 2022), the above forward and reverse process is conducted on latent embeddings $z=\mathcal{E}_{\mathrm{VAE}}(x)$, as a latent representation of raw RGB pixels $x$ encoded by a VAE (Diederik and Max, 2014; Rezende et al., 2014). And the denoised latent representation $z_{0}$ is further decoded into a normal RGB image using the VAE decoder, as $x_{0}=\mathcal{D}_{\mathrm{VAE}}(z_{0})$.
@@ B.2 ControlNet
ControlNet (Agrawala, 2023) remarkably extends the controllability of large pretrained T2I LDMs by incorporating additional conditions ($c^{f}$) like Canny edges, human pose, or depth map, which provide structural priors for the target generated image.
@@
$y_{\mathrm{c}}=\mathcal{F}(x;\Theta)+\mathcal{Z}(\mathcal{F}(x+\mathcal{Z}(c^{f};\Theta_{\mathrm{Z}1});\Theta_{\mathrm{c}});\Theta_{\mathrm{Z}2})$ (S5)
Training Objective. With the introduction of additional conditioning vector $c^{f}$, the overall training objective (Equation S2) is reformulated as:
$\mathcal{L}=\mathbb{E}_{x_{0},t,c^{t},c^{f},\epsilon\sim\mathcal{N}(0.I)}\left[\parallel\epsilon-\epsilon_{\theta}(x_{t},t,c^{t},c^{f})\parallel_{2}^{2}\right].$ (S6)
@@
Appendix C Experiments
C.1 Ablations on Layout Controls
We conduct ablation studies on individual components of the proposed layout controls, with results summarized in Tab. S1 and Fig. S1. It can be found that incorporating depth and HED controls effectively enforces structural constraints on the synthesized backgrounds, thereby enhancing semantic consistency between the foreground target and the generated scene. This structural alignment further improves the overall generation quality and visual realism. While both depth and HED controls contribute to structure-aware generation, they exhibit complementary effects: depth control produces images with clearer spatial hierarchy and depth layering, whereas HED control sharpens object boundaries and yields more distinct foreground contours. These observations highlight the necessity of jointly modeling geometric layout and edge-level structure in controllable camouflaged image generation.
C.2 Ablations on Textual-Visual Condition
We conduct ablation studies on the individual components of the proposed textual–visual condition, with results reported in Tab. S2 and Fig. S1. Simple and Detailed denote the use of a simple textual prompt and the proposed unified fine-grained textual task description, respectively, while Visual refers to the incorporation of object category information and reference background visual cues. The results show that fine-grained task descriptions consistently improve overall image generation quality. Moreover, introducing visual conditions further enhances generation fidelity by providing explicit visual guidance for background synthesis (as shown in Fig. S2), which leads to higher visual similarity and improved semantic alignment between the foreground object and the generated background.
Figure S1: Qualitative results of the ablation study on layout controls and textual-visual condition. Cst.: Contrast control. Simple: The simple textual prompt. Detailed: The unified fine-grained textual task description.
Figure S2: Illustrations of image samples retrieved by the proposed Texture-Oriented Background Retrieval (TOBR) module.
Table S1: Ablation results on layout controls. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better.
[…]
$\surd$ $\times$ $\times$ 35.29 0.0099 0.1629 0.3191 55.49 0.0146 0.2116 0.8249 87.31 0.0251 0.2007 0.7589 32.99 0.0118 0.1917 0.6343
$\surd$ $\surd$ $\times$ 32.94 0.0085 0.2627 0.3565 43.50 0.0116 0.3299 0.9049 55.91 0.0221 0.2735 0.8367 25.30 0.0104 0.2887 0.6994
$\surd$ $\times$ $\surd$ 33.76 0.0086 0.2421 0.3459 43.60 0.0127 0.3421 0.8957 56.74 0.0229 0.3006 0.8454 25.54 0.0102 0.2949 0.6957
$\surd$ $\surd$ $\surd$ 31.77 0.0071 0.3254 0.3398 35.61 0.0083 0.4027 0.8659 44.90 0.0143 0.3695 0.7938 20.30 0.0083 0.3659 0.6665
Table S2: Quantitative results of ablation study on textual-visual condition. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better. Simple: The simple textual prompt. Detailed: The unified fine-grained textual task description.
[…]
Simple $\times$ 37.59 0.0121 0.2910 0.4019 43.97 0.0151 0.3525 0.9349 61.92 0.0280 0.3160 0.8637 27.71 0.0162 0.3198 0.7335
Detail $\times$ 37.37 0.0116 0.2951 0.3639 42.94 0.0123 0.3610 0.9039 56.94 0.0235 0.3297 0.8209 26.74 0.0148 0.3286 0.6962
$\times$ $\surd$ 33.66 0.0080 0.3077 0.3591 40.92 0.0109 0.3443 0.8862 50.47 0.0191 0.3483 0.8142 23.44 0.0105 0.3334 0.6865
Detail $\surd$ 31.77 0.0071 0.3254 0.3398 35.61 0.0083 0.4027 0.8659 44.90 0.0143 0.3695 0.7938 20.30 0.0083 0.3659 0.6665
C.3 Ablations on Hyperparameters
We further investigate the impact of key hyperparameters on the inference behavior of the proposed framework. Specifically, we analyze the effects of control strength and condition strength as illustrated in Figs. S3 and S4. As these strengths increase, the generated images exhibit more saturated colors and stronger contrast, with the overall visual appearance gradually shifting from photorealism toward a painterly style.
We also examine the influence of the number of inference steps, as shown in Fig. S5. Increasing the inference steps leads to richer visual details in the generated images, at the cost of higher inference time.
Based on these observations, we adopt a balanced configuration for inference, setting the control strength to 1.0, the condition strength to 2.0, and the number of inference steps to 50, which achieves a favorable trade-off between image quality, realism, and computational efficiency.
Figure S3: Qualitative comparison of the ablation study on different control strengths. Ctrl.: Control strength.
Figure S4: Qualitative comparison of the ablation study on different condition strengths. Cond.: Condition strength.
Figure S5: Qualitative comparison of the ablation study on different inference steps. Step: Number of inference steps.
C.4 Additional Visual Results
To more comprehensively demonstrate the effectiveness of our RealCamo framework, we present additional qualitative results across three camouflage transfer scenarios, i.e., camouflaged objects transfer, salient objects transfer, and general objects transfer, as shown in Figs. S9, S10, and S11. These results highlight the superior performance and robustness of our proposed method under diverse object characteristics and transfer settings.
Table S3: Performance of MVANet (Yu et al., 2024a) on SOD and GOD with different training data. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better.
Training Data Salient Object Detection (1,000) General Object Detection (1,000)
$S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ $E_{\phi}^{a}\uparrow$ $E_{\phi}^{x}\uparrow$ $F_{\beta}^{m}\uparrow$ $F_{\beta}^{a}\uparrow$ $F_{\beta}^{x}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ $E_{\phi}^{a}\uparrow$ $E_{\phi}^{x}\uparrow$ $F_{\beta}^{m}\uparrow$ $F_{\beta}^{a}\uparrow$ $F_{\beta}^{x}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$
Original 0.911 0.933 0.934 0.943 0.885 0.876 0.902 0.869 0.033 0.745 0.783 0.833 0.835 0.656 0.683 0.688 0.626 0.086
Ori. + Gen S/G 0.901 0.926 0.923 0.939 0.873 0.858 0.897 0.849 0.039 0.730 0.767 0.815 0.816 0.635 0.675 0.670 0.606 0.095
Table S4: Performance of MVANet (Yu et al., 2024a) on the CHAMELEON and CAMO test set with different training data. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better.
Training Data CHAMELEON (76) CAMO (250)
$S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ $E_{\phi}^{a}\uparrow$ $E_{\phi}^{x}\uparrow$ $F_{\beta}^{m}\uparrow$ $F_{\beta}^{a}\uparrow$ $F_{\beta}^{x}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ $E_{\phi}^{a}\uparrow$ $E_{\phi}^{x}\uparrow$ $F_{\beta}^{m}\uparrow$ $F_{\beta}^{a}\uparrow$ $F_{\beta}^{x}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$
Original 0.915 0.929 0.957 0.953 0.871 0.867 0.895 0.846 0.022 0.870 0.896 0.918 0.927 0.842 0.847 0.865 0.809 0.055
Ori. + Gen S&G 0.905 0.927 0.911 0.971 0.846 0.811 0.904 0.796 0.031 0.826 0.842 0.881 0.885 0.760 0.784 0.795 0.717 0.079
Ori. + SynCOD12K 0.924 0.949 0.952 0.965 0.889 0.877 0.912 0.871 0.019 0.882 0.916 0.927 0.935 0.856 0.852 0.881 0.826 0.049
Table S5: Performance of MVANet (Yu et al., 2024a) on the COD10K and NC4K test set with different training data. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better.
Training Data COD10K (2,026) NC4K (4,021)
$S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ $E_{\phi}^{a}\uparrow$ $E_{\phi}^{x}\uparrow$ $F_{\beta}^{m}\uparrow$ $F_{\beta}^{a}\uparrow$ $F_{\beta}^{x}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ $E_{\phi}^{a}\uparrow$ $E_{\phi}^{x}\uparrow$ $F_{\beta}^{m}\uparrow$ $F_{\beta}^{a}\uparrow$ $F_{\beta}^{x}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$
Original 0.882 0.917 0.913 0.944 0.828 0.792 0.854 0.792 0.025 0.886 0.910 0.919 0.939 0.855 0.844 0.879 0.820 0.038
Ori. + Gen S&G 0.865 0.896 0.849 0.940 0.784 0.708 0.849 0.719 0.035 0.879 0.897 0.883 0.934 0.821 0.788 0.872 0.769 0.048
Ori. + SynCOD12K 0.896 0.934 0.919 0.953 0.850 0.805 0.877 0.818 0.021 0.891 0.919 0.914 0.938 0.860 0.840 0.885 0.824 0.037
Appendix D Downstream Applications
D.1 Model and Evaluation Metrics
Model. For downstream application evaluation, we adopt MVANet (Yu et al., 2024a) as the segmentation model. MVANet is a dichotomous image segmentation (DIS) model designed for challenging fine-grained object segmentation, typically operating on high-resolution inputs (e.g., $1024\times 1024$). Since high-resolution segmentation accuracy is not the primary focus of this work and the native output resolution of our RealCamo framework is $512\times 512$, we conduct all training, inference, and evaluation at a unified resolution of $512\times 512$ using MVANet.
Evaluation Metrics. To enable a more comprehensive evaluation, we adopt all the following metrics: mean absolute error ($M$), E-measures (including mean E-measure ($E_{\phi}^{m}$), average E-measure ($E_{\phi}^{a}$), and max E-measure ($E_{\phi}^{x}$)), F-measures (including mean F-measure ($F_{\beta}^{m}$), average F-measure ($F_{\beta}^{a}$), and max F-measure ($F_{\beta}^{x}$)), weighted F-measure ($F_{\beta}^{w}$), and structure measure ($S_{\alpha}$). Better performance is indicated by lower values of $M$ and higher values of the others.
D.2 Experiments on SOD and GOD
For the SOD task (and analogously for the GOD task), we randomly select 1,000 images from the corresponding LAKE-RED validation subset as the test set, with the remaining 5,473 images used for training. We conduct two experimental settings: (i) training the model using the original images only (denoted as “Original”), and (ii) training with both the original images and their camouflage-transferred counterparts (denoted as “Ori. + Gen S/G”). The quantitative results are reported in Tab. S3.
The results show that incorporating camouflage-transferred images as additional training data significantly degrades performance on both SOD and GOD tasks. This outcome is expected, as camouflage transfer inherently suppresses saliency cues and blurs category-specific visual characteristics, thereby obscuring the discriminative signals required for accurate detection. Consequently, models trained with such data struggle to correctly localize salient or category-specific targets, leading to inferior performance.
Figure S6: Pipeline of constructing synthetic camouflaged images in the SynCOD12K dataset.
D.3 Experiments on COD
The SynCOD12K Dataset. To enrich training data for COD, we construct a synthetic dataset termed SynCOD12K using the proposed RealCamo framework. To avoid data leakage, we start from the standard COD training split consisting of 4,040 images. For each image, the Texture-Oriented Background Retrieval (TOBR) module is employed to retrieve the top-3 background reference samples. Specifically, given an input image $\mathrm{I}$ and its retrieved reference samples $\{S_{j}\}_{j=1}^{3}$, we first extract the corresponding depth maps and HED edge maps, which are then unified to the same resolution and fused to produce a combined depth map and a combined HED map. These structural cues, together with the textual–visual conditions, are fed into the RealCamo pipeline to generate high-quality camouflaged images. The overall data generation process is illustrated in Fig. S6.
Evaluation Results. We evaluate downstream COD performance under three training configurations: (i) using the standard COD training set (denoted as “Original”), (ii) augmenting (i) with camouflaged images transferred from the salient and general subsets of LAKE-RED validation set (denoted as “Ori. + Gen S&G”), and (iii) augmenting (i) with the proposed SynCOD12K dataset (denoted as “Ori. + SynCOD12K”).
Results in Tabs. S4 and S5 show that salient- and general-synthesis camouflaged images degrade COD performance, likely due to a remaining domain gap despite camouflage transfer. In contrast, incorporating SynCOD12K consistently improves performance, indicating that realistically synthesized camouflaged images can effectively benefit downstream COD training.
Appendix E User Study
To further evaluate the quality of generated camouflaged images, we design a more comprehensive human-feedback-based evaluation framework and conduct a user study. We prepare two parallel questionnaires (i.e., Version A and Version B), each consisting of two sections. A total of 20 participants were recruited, with 10 participants assigned to each version.
We thank Wenchao Liu and Hengrui Qu for their contributions to the questionnaire design, and we also sincerely express our appreciation to the 20 participants for providing valuable human feedback results.
The questionnaires will be submitted as supplementary material, with some examples illustrated in Figs. S12 and S13.
Figure S7: Human feedback results for Part I of the user study. Top: the 1st, 2nd, and 3rd vote numbers on five questions. Bottom: the 1st and total vote percentages on five questions. Please zoom in and view in color for more details.
E.1 Part I: Multi-Dimensional Evaluation
In this part, we conduct a comprehensive multi-dimensional evaluation of camouflaged image generation, covering nine different methods inclueing DCI (Zhang et al., 2020), LCGNet (Li et al., 2022b), TFill (Zheng et al., 2022), LDM (Rombach et al., 2022), RePaint-L (Lugmayr et al., 2022), LAKE-RED (Zhao et al., 2024), CamoAny (Das and Gopalakrishnan, 2025), CT-CIG (Qian et al., 2025), and our proposed RealCamo.
This section contains 20 samples. For each sample, we provide a real camouflaged image (captured in nature) and the camouflaged targets contained in this image. We also present synthetic results generated by nine different methods. For each sample, participants are required to answer the following five questions. For each question, they should provide the top-3 most suitable answers in order of priority.
  • Q.1 Camouflage Effectiveness: Which result makes the camouflaged target hardest to detect? Consider both detection difficulty and validity of camouflage.
  • Q.2 Realism: Which result looks most like a real image? Only evaluate the nine displayed results, excluding the reference real image.
  • Q.3 Consistency: Which result is most similar to the reference real image? Consider image structure and the semantics in the content.
  • Q.4 Practicality: Which result has the largest deviation between the camouflaged target area and the ground truth? i.e., the mismatch degree between the target position in the generated result and the actual position.
  • Q.5 Comprehensive Quality: Which result has the highest image quality? Evaluate solely from the perspective of a normal image/photo.
To ensure the impartiality of the user study, the display order of the results from the nine methods is fully randomized for every example. In addition, Version A follows the sample selection protocol used in the LAKE-RED user study, whereas Version B consists of randomly selected samples.
Fig. S7 summarizes the human feedback results across the five assessment dimensions: camouflage effectiveness, realism, consistency, practicality, and comprehensive quality. RealCamo consistently achieves the highest human preference scores on all five dimensions, indicating its superior ability to generate visually coherent and convincingly camouflaged images. Notably, the most recent approaches, CamoAny and CT-CIG, also exhibit competitive performance, particularly in camouflage effectiveness and realism, suggesting that recent advances in camouflaged image generation have substantially improved perceptual fidelity.
Table S6: Human feedback results for Part II of the user study. The highest values are highlighted in bold. A lower accuracy means a higher misjudgment rate, indicating generated images with more realism.
Method Q.1 Accuracy Q.2 Confidence
Option.A Option.B Option.C Option.D
CamoAny 79.50% 51.00% 23.50% 14.00% 11.50%
CT-CIG 66.00% 49.50% 26.50% 9.00% 15.00%
RealCamo 43.00% 18.50% 20.50% 22.50% 38.50%
E.2 Part II: Further Realism Evaluation
In this part, we further evaluate the realism of the generated camouflaged images.
This section contains 30 samples. For methods with probably top-3 realism in Part I (i.e., RealCamo, CT-CIG, and CamoAny), we randomly select 10 samples for each. For each sample, we provide a real camouflaged image (captured in nature) and a camouflaged image generated by one of the abovementioned three methods. For each sample, participants are required to answer the following two questions. For each question, they should provide the most suitable answer.
  • Q.1: Which result is the real camouflaged image? i.e., the scene should naturally exist in the real world.
  • Q.2: What is your experience in making the judgment, and how confident are you in your choice? Please select one option from the following.
    • Option.A: I can instantly distinguish the real image and am fully certain of my judgment.
    • Option.B: I need a short while to make the judgment, but I am certain of my judgment.
    • Option.C: I have to deliberate carefully to make the judgment, but I am certain of my judgment.
    • Option.D: I have to deliberate carefully to make the judgment and still remain uncertain about my judgment.
To ensure the impartiality of the user study, for each sample, the method used for generating the result, and the display order of the real image are completely random.
Tab. S6 reports human feedback results on further realism evaluation, including the judgment accuracy and confidence. RealCamo induces the highest misjudgment rate (i.e., the lowest accuracy), requires a longer average decision time, and elicits the greatest uncertainty among participants. These results demonstrate that images produced by RealCamo are perceptually challenging even for human observers, which further validates the effectiveness of our proposed framework.
Appendix F Limitations and Future Work
Despite the superior performance of RealCamo in realistic camouflaged image synthesis and its demonstrated benefits for downstream COD tasks, several limitations remain and motivate future research.
First, camouflage transfer from salient and general images remains challenging, as shown in Fig. S8 (a). Although RealCamo achieves reasonable transfer results, the synthesized images transferred from salient and general images still exhibit weaker visual camouflage. This may stem from both the domain gap between naturally camouflaged training data and non-camouflaged source images, as well as the interaction between explicit structural controls and pretrained diffusion priors, which can impede the emergence of strong camouflage cues. Developing mechanisms that enhance visual camouflage while preserving structural consistency is an important direction.
Second, generation diversity is inherently constrained by structural controls. While such controls improve realism and semantic consistency, they limit variability in the generated scenes. Although our controlled information fusion (as shown in Fig. S6) partially mitigates this issue, designing principled methods that jointly support structural guidance and diverse generation remains an open problem.
Third, image resolution and fine-grained details are limited by the underlying Stable Diffusion 1.5 backbone, which operates at a $512\times 512$ resolution. Compared to high-resolution real camouflaged images, the generated results lack fine details, as demonstrated in Fig. S8 (b). Leveraging more advanced high-resolution generative models (e.g., SDXL (Podell et al., 2024) or FLUX (Labs et al., 2025)) is a promising avenue for improving visual fidelity.
Finally, while we show that synthesized camouflaged images can benefit downstream COD training, how to more effectively integrate synthetic data into COD remains underexplored and warrants further investigation, with some pioneer work (Luo et al., 2025; Chen et al., 2026) here for reference.
Figure S8: Limitations of the proposed RealCamo framework. (a) Camouflage transfer from salient and general images remains challenging. (b) Compared to high-resolution real camouflaged images, the generated lower-resolution result lacks fine details.
Figure S9: Additional qualitative comparison with existing SOTA methods on camouflaged objects transfer. Our results are demonstrated in the rightmost column with red box.
Figure S10: Additional qualitative comparison with existing SOTA methods on salient objects transfer. Our results are demonstrated in the rightmost column with red box.
Figure S11: Additional qualitative comparison with existing SOTA methods on general objects transfer. Our results are demonstrated in the rightmost column with red box.
Figure S12: An example for Part I of the user study (Version A).
Figure S13: Some examples for Part II of the user study (Version A).